OpenStudy (anonymous):

can someone help me. i'm not getting anything done on problem set1. i dont wanna give it up. just not getting any code work.

OpenStudy (e.mccormick):

Well, what sort of errors are you having? Where it is being a problem?

OpenStudy (anonymous):

e.mccormick @e.mccormick and others K! so far i have been able to travel to this extend: #Initial balance =IB # Annual interest rate =AIR #monthly interest rate = MIR #Annual interest =AI #minimum monthly payment rate =MMPR #2% #Mininmum monthly payment = MMP #monthly interest rate =MIR #principal paid #principal paid #remaining balance #Remaining balance =RB IB = float (raw_input('enter the outstanding balance on credit card:')) AIR = float(raw_input ('enter the annual credit card interest rate as a decimal:')) MMPR = float (raw_input ('enter minimum monthly paymment rate in decimal:')) MMP = MMPR*IB MIR = AIR/12.0 MR = MIR*IB PP = MMP- MR RB = IB - PP while..... my problem seem to set in here where i need to introduce while loop to run the subsequent months balances. anyhelp

OpenStudy (e.mccormick):

You look at what things would need to be recalculated every month and put those inside the loop. For example, every month there is interest. Adding in the interest needs to be inside the loop.